one root of quadratic equation 3x square - 10 x +3=o is 1/3.find other root
Found 2 solutions by Grinnell, stanbon:
Answer by Grinnell(63)

In factoring we break the polynomial down and then 'hit or miss'
3x^2-10x+3=0
(3x-?)(x-??) = 0
we know that product of (?) and (??) equals positive three. (-something times -something=positive something!)

how about for starters one and three?
let 1=?, let 3 =??, substitute--this works. THE OTHER ROOT IS THREE WHICH IS THE ANSWER!
(understand for the first root 3x-1 = 0, yes x equals 1/3.
We solved for the other.

Answer by stanbon(75887)

one root of quadratic equation 3x square - 10 x +3=o is 1/3.find other root
-----
If 1/3 is a root, (3x-1) is a factor:
----
Use Snythetic division to find the quotient:
1/3)....3....-10....3
........3......-9..|..0
Quotient: 3x-9
---
Solve 3x-9 = 0
x = 3
-------------------
The other root is x = 3
==============================
Cheers,
Stan H.